# NAMI: Efficient Image Generation via Progressive Rectified Flow Transformers

### v1

@@ 4.3 Ablation Study
Performance of NAMI Components In addition to generation quality, we evaluated the impact of different components of the NAMI architecture on inference time. We performed 30 inference steps on an 80GB VRAM A100 GPU, excluding the time spent on the VAE [21] and text encoders. We present in Figure 1 the variation in inference advantage of our method as the resolution increases. We also evaluated the inference time advantages brought by adjusting the flow and transformer components, respectively.
As shown in Figure 9, the FLUX-based method at a resolution of 1024 shows a 28.68% reduction in inference time when only flow piecewise is applied. Additionally, with model partitioning, the inference time further decreased by 11.81%.
Figure 9: Ablation study on the inference time of NAMI.
@@ 5 Conclusion
In this paper, we introduce a novel multi-stage flow and spatially nested model architecture and employs a multi-resolution joint training strategy to accelerate model convergence while ensuring performance. By dividing the rectified flow into multiple stages based on resolution and progressively increasing the number of transformer layers as the resolution grows, our method efficiently balances computational cost and performance. The proposed NAMI architecture reduces inference time by 40% at 1024 resolution. Additionally, we have constructed the NAMI-1K benchmark to address the distribution biases and data leakage issues present in open-source benchmarks. The results demonstrate that our model is competitive with current state-of-the-art methods.

### v2

@@ 5 Conclusion
In this paper, we introduce a bridge-connected multi-stage flow and spatially nested model architecture and employs a multi-resolution joint training strategy to accelerate model convergence while ensuring performance. By dividing the rectified flow into multiple stages based on resolution and progressively increasing the number of transformer layers as the resolution grows, our method efficiently balances computational cost and performance. The proposed NAMI architecture reduces inference time by 64% at 1024 resolution. Additionally, we have constructed the NAMI-1K benchmark to address the distribution biases and limited prompt diversity issues present in open-source benchmarks. The results demonstrate that our model is competitive with current state-of-the-art methods.
@@ B. More Results of NAMI
C. The inference time of NAMI Components
We performed 30 inference steps on an 80GB VRAM A100 GPU, excluding the time spent on the VAE and text encoders. We evaluated the inference time advantages brought by adjusting the flow and transformer components, respectively. As shown in Figure 25, the FLUX-based method at a resolution of 1024 shows a 53.01% reduction in inference time when only flow piecewise is applied. Additionally, with model partitioning, the inference time further decreased by 11.81%.
Figure 25: The inference time of NAMI Components.
D. The visualization of different methods